81. Suppose cos(6) and is in quadrant IV. Find sin(0) and tan , (X) 82. Evaluate: co 83. Evaluate: sin(-600°) question 81 show all work please
Answer #1

For Powth quedrant cosa = Posative Since tance = negative given cosce=? , ato find Sino and tan o, first find opposite Siclehypot nuse = base + Altitude? 32-22+ Altrade? Altitude?=9-455 Altitude = 15 so opposite side will be V5 il as lo Now find sin